OpenStudy (anonymous):

Use Lagrange multipliers to solve the given optimization problem. Find the maximum value of f(x, y) = xy subject to x + 2y = 52. @whpalmer4

OpenStudy (whpalmer4):

Sure, now I have to learn about Lagrange multipliers, eh? :-)

OpenStudy (anonymous):

haha i just keep testing you!

OpenStudy (fibonaccichick666):

calc 3?

OpenStudy (anonymous):

idk if its labeled as calc 3 but calculus yeah!

OpenStudy (fibonaccichick666):

ahh ok

ganeshie8 (ganeshie8):

say, g is the constraint function, then the min/max value occurs when : \(\bigtriangledown f = \lambda \bigtriangledown g\) find the partials fx, fy, gx, gy and setup below equations : \(f_x = \lambda g_x\) \(f_y = \lambda g_y\) and solve

ganeshie8 (ganeshie8):

u able to find \(\lambda\), x, y values?

ganeshie8 (ganeshie8):

wid the constraint equation, u have 3 equations and 3 unknowns to find.. piece of cake to solve

OpenStudy (ybarrap):

$$ f(x, y) = xy\\ g(x,y)= x + 2y = 52\\ \Lambda(\lambda,x,y)=f(x,y)-\lambda(g(x,y)-c)\\ \Lambda(\lambda,x,y)=xy-\lambda(x+2y-52)\\ \Lambda_x=y-\lambda=0\\ \Lambda_y=x-2\lambda=0\\ \Lambda_{\lambda}=x+2y-52=0\\ $$ Solve for x and y.

OpenStudy (ybarrap):

$$ y=\lambda\\ x-2y=0\\ x=2y\\ 2y+2y-52=0\\ 4y=52\\ y=\cfrac{52}{4}\\ $$ That's it! Any questions?

OpenStudy (anonymous):

no that all makes sense!! thank you :)

OpenStudy (ybarrap):

yw